A company hired Ronnie and Kenny as consultants. Together they made 14,500 dollars. If Ronnie had earned 500 dollars less, both he and Kenny would have been paid the same. How much were they each paid?
Answer by RAY100(1637)

Let "R" = Ronnie pay,, and Let "K" = Kennie pay
.
(1) R + K = 14,500
.
(2) (R-500) = K
.
substitute (2) in (1)
.
R + (R-500) = 14,500
.
2R = 15,000
.
R = $ 7500
.
substituting in (1)
.
7500 = K = 14,500
.
K = $ 7,000
.
checking
(1) 7500 + 7000 = 14500 ,,,,,ok
.
(2) ( 7500 -500) = 7000,,,,,ok